If any one can help me I will appreciate it very much. This is mu last chance at understanding befrore my final tomorrwo. pLESE HELP ME SOLVE x^2 -5x +6 =6, AND 3a^2 + 14a + 8 =0 Ipromise this will be the last time tonight. Thank you for any help.
Answer by venugopalramana(3286)

If any one can help me I will appreciate it very much. This is mu last chance at understanding befrore my final tomorrwo. pLESE HELP ME SOLVE x^2 -5x +6 =6,
X^2-5X=0
X(X-5)=0
X=0 OR X=5
AND 3a^2 + 14a + 8 =0
3A^2+12A+2A+8=0
3A(A+4)+2(A+4)=0
(3A+2)(A+4)=0
3A+2=0...OR...A=-2/3
A+4=0.....OR...A=-4

Ipromise this will be the last time tonight. Thank you for any help.